# Post-Newtonian Tests of Gravitational Quantum Field Theory with Spin and Scaling Gauge Symmetry

### v1

@@ Abstract
keywords:
Quantum gravity , Quantum field theory , Post-Newtonian limit , Gravitation
journal: Physics Letters B
1 Introduction
@@
2 PN metric solution of GQFT
2.1 Field equation
@@ 2.2 PN expansions
$\Gamma^{\mu}_{\nu\lambda}=\Gamma^{\mu}_{\nu\lambda(2)}+\Gamma^{\mu}_{\nu\lambda(4)}+...,$
@@
$\Gamma^{\mu}_{\nu\lambda}=\Gamma^{\mu}_{\nu\lambda(3)}+....$
@@ 2.3 PN metric solution
3 Experimental constraints on GQFT
@@
From the above discussions, the experimental constraints on the key Eddington-Robertson-Schiff parameters $\gamma$ and $\beta$ will further set bounds on the combined coupling parameter $\gamma_{W}$ of GQFT in the weak field and slow motion limit. The strongest bound is from the studies of anomalous perihelion shift of Mercury’s orbit, which gives $|\gamma_{W}|=|\gamma_{G}(\alpha_{G}-\alpha_{W}/2)|\leq(2.4\pm 30)\times 10^{-6}$, which improves a former constraint $|\gamma_{W}\leq\mathcal{O}(10^{-5})|$ [43]. Such a bound implies that the mass scale of spin gauge field should be around or below the energy scale of grand unification theory. Moreover, one can further conclude that, under such bound, the new GQFT had successfully passed today’s most stringent and keynote tests in experimental relativity, including precision measurements of light deflection, Shapiro time delay, perihelion shift of Mercury’s orbit, Lense-Thirring effect of satellite orbit, frame-dragging precession of orbiting gyroscope etc.
4 Concluding remarks

### v2

@@ Abstract
I Introduction
@@
II PN metric solution of GQFT
II.1 Field equation
@@
It is worth to notice that the GQFT model is related to the so called teleparallel approach to gravity, in which the localized translation symmetry plays the crucial role and the tetrad field arises as the fundamental field variable, please see [28, 29] for details. In a recent work [30], the PPN expansion of the general formalism of teleparallel gravity was worked out and the corresponding constraints from observations were obtained, and the comparison between these two models is discussed in the next section.
@@ II.2 PN expansions
$\Gamma^{\mu}_{\nu\lambda}=\Gamma^{\mu}_{\nu\lambda(2)}+\Gamma^{\mu}_{\nu\lambda(4)}+...,$ (7)
@@
$\Gamma^{\mu}_{\nu\lambda}=\Gamma^{\mu}_{\nu\lambda(3)}+....$ (8)
@@ II.3 PN metric solution
III Experimental constraints on GQFT
@@
From the above discussions, the experimental constraints on the key Eddington-Robertson-Schiff parameters $\gamma$ and $\beta$ will further set bounds on the combined coupling parameter $\gamma_{W}$ of GQFT in the weak field and slow motion limit. The strongest bound is from the studies of anomalous perihelion shift of Mercury’s orbit, which gives $|\gamma_{W}|=|\gamma_{G}(\alpha_{G}-\alpha_{W}/2)|\leq(2.4\pm 30)\times 10^{-6}$, which improves a former constraint $|\gamma_{W}\leq\mathcal{O}(10^{-5})|$ [46]. Such a bound implies that the mass scale of spin gauge field should be around or below the energy scale of grand unification theory. Moreover, one can further conclude that, under such bound, the new GQFT had successfully passed today’s most stringent and keynote tests in experimental relativity, including precision measurements of light deflection, Shapiro time delay, perihelion shift of Mercury’s orbit, Lense-Thirring effect of satellite orbit, frame-dragging precession of orbiting gyroscope etc. For the PPN approximation of a general teleparallel gravity theory, it is found that [30] the theory does not exhibit any violations in Einstein’s equivalence principle and conservation law of energy-momentum. It contains only two nontrivial PN parameters $\gamma$ and $\beta$, and, in some sense similar to our results in the last section, the deviations in these parameters from GR depend only on a unique parameter, see again [30] for detailed discussions. This manifests, in some sense, the potential links between the GQFT model considered in this work and general teleparallel gravity.
IV Concluding remarks